# Relightful Harmonization: Lighting-aware Portrait Background Replacement

### 11 Dec 2023 v1

@@ 1 Introduction
There are two principal sets of methods that could automatically adjust the foreground to match the background for human portraits: (1) image harmonization techniques, and (2) portrait relighting methods. Harmonization-based methods [7, 5, 56, 14, 26, 8, 15, 16, 22, 29, 62] aim to match the color statistics of the foreground with those of the background, by adjusting the foreground color tone, contrast, and illumination. Yet, they often overlook the lighting characteristics and leave the foreground illumination effects unchanged, such as the lighting direction and shadows, potentially resulting in an unnatural appearance when the background has distinct lighting conditions. For instance, compositing a person photographed under a top-down light into a sunset scene might make the composite non-realistic to the human eye. On the other hand, recent work on portrait relighting [36, 40, 51, 73, 70, 66, 59, 63, 32] are designed for matching the lighting of the subject towards a new environment by using the paired training data acquired with the light stage system [10]. Nevertheless, current relighting methods often require dynamic range (HDR) panorama environment maps [40] during training and inference, which are not always feasible to acquire, especially in casual photography settings. In this work, we explore the possibility of generating realistic and lighting-aware composition images in a straightforward harmonization set up. Given a foreground image (with its corresponding alpha mask) and an arbitrary background image, we propose a unified and end-to-end framework that encompasses both color and lighting harmonization. We approach the task through a conditional generative framework, leveraging a pretrained diffusion model [8, 3], and develop a three-stage training pipeline.
@@ 2 Related Works
Image Harmonization aims to rectify color, contrast, and style differences between foreground and background to ensure composite images look natural and consistent. In deep learning, this task is approached as an end-to-end image-to-image translation problem [74, 52, 16, 22, 7, 15, 8, 56, 5, 14], where the network is trained to predict a harmonized image from the input composite. Pixel-aligned datasets are usually created by altering foreground color in real images with pre-designed [7] or learned [39] augmentations.
@@
Nevertheless, existing harmonization methods primarily focus on global color adjustment, and overlook the subtle but important discrepancies in foreground and background lighting, i.e., direction, intensity, and shadow effects. This oversight can lead to a harmonized image that, despite matching colors, still appears unnatural due to mismatched lighting conditions. Therefore, we postulate that enhancing the lighting-awareness of harmonization models is a vital yet underexplored area for obtaining natural and realistic composition.
@@
Portrait Relighting Recent advancements in portrait relighting has been driven by deep learning methods, [36, 40, 51, 73, 70, 66, 59, 63, 32, 42]. They leverage supervised training with the paired training data acquired with the light stage system [10]. These methods require a target HDR enviroment map as the external input source, and typically involve the intermediate prediction of surface normals, albedo, and/or a set of diffuse and specular maps with ground truth supervision. However, reliance on HDR maps for background replacement and harmonization tasks poses significant limitations on their applicability in everyday scenarios where HDR maps cannot be easily captured alongside [40]. In our framework, we do not require any external input sources except a single target background image. Furthermore, many current relighting systems employ multistage frameworks and/or heavily rely on external packages [12]. The accuracy and performance of these systems are consequently contingent on the precision of each individual stage, making the overall process complicated and prone to errors propagated through these intermediate steps. Additionally, the datasets commonly employed for training are rendered from limited light stage illumination acquisition, which means the target images utilized during the training phase are not captured in real-world conditions but are rendered composites. Therefore, its generalization on unseen images, as well as the photorealism when applied to arbitrary background replacement tasks remain unclear.
@@
Diffusion Models [49, 19, 50, 11] have significantly advanced the image and video synthesis quality [43, 45, 18, 24, 20, 25]. Image-conditioned diffusion models [67, 57, 46] typically take an image as additional input to perform an image-conditioned generation such as image enhancement [28, 23, 44, 25], harmonization [6, 31, 30] and translation [67, 48, 72, 57, 27, 38], typically trained on task-specific pairwise data. Recently, the application of pretrained text-to-image diffusion models [37, 45, 47] has been extended to image editing tasks [33, 17, 9, 3, 35, 55, 2, 53, 34, 71, 60, 21, 4, 1, 13]. These models leverage text-image correlations to perform context modifications in image editing, such as ‘adding a sunset’ with InstructPix2Pix [3], which is loosely connected to our lighting-aware set up. However, text-based editing does not incorporate spatial information, thus lacking finegrained control to the model. Instead, we propose to use a spatial lighting representation as the new ‘Instruction’ that guide the diffusion model to perform lighting-aware editing.
@@ 3 Method
Stage III: Finetuning for Photorealism: In the third stage, we finetune our model using high-quality pixel-aligned training pairs from natural images, where these paired datasets are generated via a novel data synthesis pipeline using the stage 2 model as a data augmenter, as shown in Fig. 2-III.
@@ 3.1 Lighting-aware Diffusion Training
Lighting Conditioning: We further modify the diffusion backbone to incorporate explicit lighting conditions. As depicted in Fig. 2-I, a lighting conditioning branch is integrated atop the UNet backbone, injecting a lighting representation encoded from the target background image by a convolutional neural network (CNN) , at multiple scales within the UNet. The conditioning mechanism is designed in a similar fashion as ControlNet [67] where conditional feature maps are added to the UNet features at respective resolution within the encoder. Details of the network architecture can be found in the appendix.
With the light stage dataset, we postulate that the lighting representation can be learned from the pairwise training. Specifically, a training tuple from the same subject is sampled as . Noise is progressively added to the target image until time step , resulting in a noisy image . The UNet, denoted as , is conditioned on the background-extracted lighting feature , and is trained to predict the noise , with the following diffusion model objective:
@@ 3.2 Lighting Representation Alignment
Then, we freeze the environment-conditioned model and introduce an alignment network that calibrates the background-extracted lighting representation to align with its environment map equivalent. We formulate such a process as an inverse problem that can be learned with a network , under a supervised loss. For a training tuple , the alignment network takes as input, and maps it with the alignment network to . The environment extracted feature is utilized as the ground truth, and we use a objective for this stage defined as follows:
During this phase, we update only and , while freezing the other networks. We assume that this alignment will enable the background-derived lighting representation to more accurately encode the environmental lighting, which is empirically verified in our ablation as the aligned feature maps reflect more global illumination information shown in Fig. 7. Once trained, we integrate the aligned feature extraction and conditioning branch into the environment-conditioned UNet , formulating our final model architecture as in Fig. 2-III.
@@ 3.3 Finetuning for Photorealism
The light stage dataset serves as a valuable resource for learning lighting representations, providing physically constrainted relighting pairs. However, it is essential to recognize that this dataset is compositional by nature. It uses backgrounds projected from environment maps, combined with relit foreground subjects, to create composites that serve as ground truth for the diffusion model. However, these composites differ from real photographs, leading to potential concerns about the photorealism of the model’s outputs. Moreover, due to the cost of light stage data acquisition, it restricts the number and diversity of the subjects that can be collected. The diversity of the background images is also bounded by the scale of available environment maps during rendering (usually a few thousands). These limitations could affect the model’s generalization ability to real images, as well as the capacity to produce realistic and varied lighting effects.
To this end, we propose a third stage that finetunes our final model for improved photorealism, outlined in Fig. 2 III.
@@ 4.1 Setup and Metrics
To evaluate the proposed method, three testing scenarios are created: (1) 500 Light stage rendered test pairs to evaluate the lighting accuracy; (2) 200 natural image test pairs, synthetically created from real images using our data synthesis pipeline, to assess the lighting realism; and most importantly, (3) Real-world portraits combined with arbitrary backgrounds, examining the model’s generalizability and adaptability in real-life scenarios. For (1) and (2), we also quantify the results with common metrics MSE, SSIM, PSNR and LPIPS [68].
To benchmark, we compare Relightful Harmonization with both established harmonization methods INR [5], PCT [14], Harmonizer [26] and PIH [56], and relighting method TR [40]. We also construct a relighting baseline with a transformer architecture and trained it with our light stage data. More details are provided in the appendix. Note that TR [40] is applied only on the light stage test set as they are not applicable without HDR maps.
@@ 4.2 Implementation Details
Our model is implemented in PyTorch [41] using GB A100 at resolutions, with 96 batch size. In stage I, we initialize UNet from the pretrained weights of the InstructPix2Pix [3] checkpoint. In the first and second stage, we use in total 400 training image pairs rendered with light stage data. We set learning rate to . In the third stage, we train the network with additional 200 pairs of images synthesized from natural images. More details are provided in the appendix.
@@ 4.4 Arbitrary Portrait Background Replacement
We further explore our framework on various in-the-wild portraits for compositing and harmonizing them with arbitrary natural backgrounds.
@@
Creative Background Replacement: Our method allows for creative background replacement for portrait images. In Fig. 4(d) (right), we create a sequence of background crops from one panorama image with a sliding window, so that the major light source consistently changes from left to right. We observe visually consistent lighting changes on the foreground subject. Similarly, Fig. 4(d) (left) shows the harmonization results by placing a subject into different timelapse video frames at different timepoints, resulting in a series of portrait images that effectively mimicked a timelapse effect.
Reference-based Harmonization: Our pipeline also extends to reference-based harmonization, where users can integrate their portrait photos with scenes from a chosen reference portrait image. This is achieved by first inpainting the reference portrait to generate a background image, onto which the desired foreground is composited. Our method ensures that the final output matches the tone and lighting of the original scene. Fig. 4(e) showcases examples where our results effectively align with the tone and lighting of the reference image, displayed in the top left.
@@ 5 Conclusion
We present Relightful Harmonization, a novel lighting-aware diffusion model tailored to blend advanced lighting effects into foreground portraits when compositing onto diverse background images. Our method surpasses existing benchmarks, showing superior performance in real-world scenarios. However, limitations exist including a resolution cap of 512x512, potentially affecting facial detail, especially in smaller faces. Also, subtle variations may occur in the subject’s clothing and skin tones. Detailed analyses and additional failure cases are included in the appendix.

### 07 Apr 2024 v2

@@ 1 Introduction
There are two principal sets of methods that could automatically adjust the foreground to match the background for human portraits: (1) image harmonization techniques, and (2) portrait relighting methods. Harmonization-based methods [7, 5, 62, 14, 27, 8, 15, 16, 23, 30, 69, 57, 68, 31] aim to match the color statistics of the foreground with those of the background, by adjusting the foreground color tone, contrast, and illumination. Yet, they often overlook the lighting characteristics and leave the foreground illumination effects unchanged, such as the lighting direction and shadows, potentially resulting in an unnatural appearance when the background has distinct lighting conditions. For instance, compositing a person photographed under a top-down light into a sunset scene might make the composite non-realistic to the human eye. On the other hand, recent work on portrait relighting [39, 43, 56, 80, 77, 73, 65, 70, 34] are designed for matching the lighting of the subject towards a new environment by using the paired training data acquired with the light stage system [10]. Nevertheless, current relighting methods often require dynamic range (HDR) panorama environment maps [43] during training and inference, which are not always feasible to acquire, especially in casual photography settings. In this work, we explore the possibility of generating realistic and lighting-aware composition images in a straightforward harmonization set up. Given a foreground image (with its corresponding alpha mask) and an arbitrary background image, we propose a unified and end-to-end framework that encompasses both color and lighting harmonization. We approach the task through a conditional generative framework, leveraging a pretrained diffusion model [8, 3], and develop a three-stage training pipeline.
@@ 2 Related Works
Image Harmonization aims to rectify color, contrast, and style differences between foreground and background to ensure natural and consistent composition. In deep learning, this task is approached as an end-to-end image-to-image translation problem [81, 58, 16, 23, 7, 15, 8, 62, 5, 14, 57, 68, 31], where the network is trained to predict a harmonized image from the input composite. Pixel-aligned datasets are created by altering foreground color in real images with pre-designed [7] or learned [42] augmentations.
@@
Yet, existing harmonization methods primarily focus on global color adjustment, overlooking the subtle but important discrepancies between foreground and background lighting, i.e., direction, intensity, and shadow. This can lead to a harmonized image that, despite matching colors, still appears unnatural due to mismatched lighting conditions. Therefore, we postulate that enhancing the lighting-awareness of harmonization models is a vital yet underexplored area for natural and realistic composition.
@@
Portrait Relighting Recent advancements in portrait relighting has been driven by deep learning methods [39, 43, 56, 80, 77, 73, 65, 70, 34, 45, 35]. They leverage supervised training with the paired training data acquired with the light stage system [10]. These methods require a target HDR enviroment map as the external input source, and typically involve the intermediate prediction of surface normals, albedo, and/or a set of diffuse and specular maps with ground truth supervision. However, reliance on HDR maps for background replacement and harmonization tasks poses significant limitations on their applicability in everyday scenarios where HDR maps cannot be easily captured alongside [43]. In our framework, we do not require any external input sources except a single target background image. Furthermore, many current relighting systems employ multistage frameworks and/or heavily rely on external packages [12]. The accuracy and performance of these systems are consequently contingent on the precision of each individual stage, making the overall process complicated and prone to errors propagated through these intermediate steps. Additionally, the datasets commonly employed for training are rendered from limited light stage illumination acquisition, which means the target images utilized during the training phase are not captured in real-world conditions but are rendered composites. Therefore, its generalization on unseen images, as well as the photorealism when applied to arbitrary background replacement tasks remain unclear.
@@
Diffusion Models [52, 20, 53, 11] have significantly advanced the image and video synthesis quality [46, 48, 19, 25, 21, 26]. Image-conditioned diffusion models [74, 63, 49] typically take an image as additional input to perform an image-conditioned generation such as image enhancement [29, 24, 47, 26], harmonization [6, 33, 32] and translation [74, 51, 79, 63, 28, 41], typically trained on task-specific pairwise data. Recently, the application of pretrained text-to-image diffusion models [40, 48, 50] has been extended to image editing tasks [36, 18, 9, 3, 38, 61, 2, 59, 37, 78, 66, 22, 4, 1, 13, 54, 55, 17]. These models leverage text-image correlations to perform context modifications in image editing, such as ‘adding a sunset’ with InstructPix2Pix [3], which is loosely connected to our lighting-aware set up. However, text-based editing does not incorporate spatial information, thus lacking finegrained control to the model. Instead, we propose to use a spatial lighting representation as the new ‘instruction’ that guide the diffusion model to perform lighting-aware editing.
@@ 3 Method
Stage III: Finetuning for Photorealism: In the third stage, we finetune our model using high-quality pixel-aligned training pairs from natural images, where these paired datasets are generated via a novel data synthesis pipeline (Fig. 2-III) using the stage 2 model as a data augmenter.
@@ 3.1 Lighting-aware Diffusion Training
Lighting Conditioning: We further modify the diffusion backbone to incorporate explicit lighting conditions. As depicted in Fig. 2-I, a lighting conditioning branch is integrated atop the UNet backbone, injecting a lighting representation encoded from the target background image by a CNN , at multiple scales within the UNet. The conditioning mechanism is designed in a similar fashion as [74] where conditional feature maps are added to the UNet features at respective resolution within the encoder.
With the light stage dataset, we postulate that the lighting representation can be learned from the pairwise training. Specifically, a training tuple from the same subject is sampled as . Noise is progressively added to the target image until time step , resulting in a noisy image . The UNet, denoted as , is conditioned on the background-extracted lighting feature , and is trained to predict the noise , with the following objective:
@@ 3.2 Lighting Representation Alignment
Then, we freeze the environment-conditioned model and introduce an alignment network that calibrates the background-extracted lighting representation to align with its environment map equivalent. We formulate such a process as an inverse problem that can be learned with a network , under a supervised loss. For a training tuple , the alignment network takes as input, and maps it with the alignment network to . The environment extracted feature is utilized as the ground truth, and we use a objective:
During this phase, we update only , while freezing the other networks. We assume that this alignment enhances the background-derived lighting representation to more accurately encode the environmental lighting, which is empirically verified where the aligned feature maps reflect more global illumination information (Fig. 7). Once trained, we integrate the aligned feature extraction and conditioning into , formulating our final model in Fig. 2-III.
@@ 3.3 Finetuning for Photorealism
The light stage dataset serves as a valuable resource for learning lighting representations, providing physically constrainted relighting pairs. However, it is essential to recognize that this dataset is compositional by nature. It uses backgrounds projected from environment maps, combined with relit foreground subjects, to create composites that serve as ground truth for the diffusion model. However, these composites differ from real photographs, leading to potential concerns about the photorealism of the model’s outputs. Moreover, due to the cost of light stage data acquisition, it restricts the number and diversity of the subjects that can be collected. The diversity of the background images is also bounded by the scale of available environment maps during rendering (a few thousands). These limitations could affect the model’s generalization ability to real images, and the capacity to produce realistic and varied lighting effects.
Therefore, we propose a third stage (Fig. 2 III) that finetunes our final model for improved photorealism.
@@ 4.1 Setup and Metrics
Three testing scenarios are created for evaluation: (1) 500 Light stage rendered test pairs to evaluate the lighting accuracy; (2) 200 natural image test pairs, synthetically created from real images using our data synthesis pipeline, to assess the lighting realism; and most importantly, (3) Real-world portraits combined with arbitrary backgrounds, examining the model’s generalizability and adaptability in real-life scenarios. For (1) and (2), we also quantify the results with common metrics MSE, SSIM, PSNR and LPIPS [75]. To benchmark, we compare Relightful Harmonization with both established harmonization methods INR [5], PCT [14], Harmonizer [27] and PIH [62], and relighting method TR [43]. We also construct a relighting baseline with a transformer architecture and trained it with light stage data. More details are provided in the appendix. Note that relighting methods are applied only on the light stage test set as they are not applicable without HDR maps.
@@ 4.2 Implementation Details
Our model is implemented in PyTorch [44] using GB A100 at resolutions, with 96 batch size. In stage I, we initialize UNet from the pretrained weights of the InstructPix2Pix [3] checkpoint. In the first and second stage, we use in total 400 training image pairs, rendered from a arbitrary combination of 100 unique light stage subjects, and 2908 HDR environment maps. We also randomly rotate the HDR maps and use various FoVs to increase the diversity of the background. In the third stage, we train the network with additional 200 pairs of images synthesized from natural images. We set learning rate to . More details are provided in the appendix.
@@ 4.4 Arbitrary Portrait Background Replacement
We further test our model on various in-the-wild portraits by compositing them onto arbitrary natural backgrounds.
@@
Creative Background Replacement:
Our method allows for creative background replacement for portrait images. In Fig. 4(d) (right), we create a sequence of background crops from one panorama image with a sliding window, so that the major light source consistently changes from left to right. We observe visually consistent lighting changes on the foreground. Similarly, Fig. 4(d) (left) shows the harmonization results by placing a subject into different timelapse video frames at different timepoints, resulting in a series of portrait images that effectively mimicked a timelapse effect.
Reference-based Harmonization:
Our pipeline also extends to reference-based harmonization, where users can integrate their portrait photos with scenes from a chosen reference portrait image. This is achieved by first inpainting the reference portrait to generate a background image, onto which the desired foreground is composited. Our method ensures that the final output matches the tone and lighting of the original scene. Fig. 4(e) showcases examples where our results effectively align with the tone and lighting of the reference image, displayed on the top left.
@@ 5 Conclusion
We present Relightful Harmonization, a novel lighting-aware diffusion model to blend advanced lighting effects into foreground portraits when compositing onto diverse background images. Limitations exist including a resolution cap of 512x512, potentially affecting facial detail, especially in smaller faces. Also, subtle variations may occur in the subject’s clothing and skin tones. Detailed analyses and additional failure cases are included in the appendix.
6 Acknowledgement
We are grateful for Yannick Hold-Geoffroy who provides the panorama environment maps and set up the rendering scripts. We thank Chaowei Company for the support of light stage dataset. We thank David Futschik for running the testing on Total Relighting. We thank Scott Cohen for the inspiration for our project name.